Tuesday, July 31, 2012

Glycol System & Filter Modules

Today's larger achievements include Mark and Charles'  test run to install filter modules (with aluminum blanks instead of real glass!) into the instrument. The first two filters are scheduled for installation this Friday, with the remaining two filters following early next week.

Before the filters will be installed the glycol cooling system will be fully tested. Today the glycol system was filled up, and air is bleeding out right now. The instrument heat exchangers will see glycol tomorrow, but they have been already successfully tested in Tucson. 

While everybody else had fun up therein the dome, Andrey I spent most of the day in the control room configuring user IDs to work well with the network file system and writing shutdown procedures and alike. Doen't make good pictures, but it is important since we are approaching the time when Kitt Peak personell will start to look after ODI while the team is not on the mountain. 

Tomorrow we will finish up the instrument connections and attach the power supply box. Thursday we should be able to operate detectors for the first time on the mountain. 

Daniel




Charles inspects the filter arm.

A first filter module  is installed in ODI. The yellow protective covers and lifting fixture are still attached. 

Flushing the glycol cooling system. As a precaution, ODI is covered with plastic sheeting.

One more video

Here is the last of the three installation time lapse movies. It shows the installation of the dewar and the cabling the following days. It is raw and uncut again since Pete will do a more comprehensive editing later. Enjoy.

Daniel



Installation Part III from Daniel Harbeck on Vimeo.

Monday, July 30, 2012

Still plugging along

The last few days have not been as picturesque as the installation of the large ODI parts. Mark Hunten is busy plugging in all cable, fiber, and to replace one last cable for one of the two focus CCDs. Tomorrow we plan to go ahead and flush out the glycol cooling system – note that we have to remove about 1.2 kW of heat from the CCD controller boxes.

Once we can be sure that there are no leaks in the glycol system, we will start to cool down the focal plane and install the first of the filters later this week. Amazingly, so far everything is proceeding pretty much as planned.

In the meantime, enjoy the first iteration of time lapse movies of the ODI installation. The movies below show the installation of of forward corrector and the instrument support package.  Once Pete will return from vacation we will post his director's cut here, as well.

Daniel



pODI Installation Part I from Daniel Harbeck on Vimeo.



pODI Installation Part II from Daniel Harbeck on Vimeo.

Friday, July 27, 2012

Insert tab A in ... which one is slot A?

Mark, Daniel, and Charles spent yesterday and today connecting (most of) the wires, fibers, hoses, cables, and tubes on ODI.  They have one or two more days before we turn the whole instrument on.  Everything is going well - other than the discovery that the chain that holds the cable (in an elongated loop in the bottom photo) is not the right size.  A replacement has been ordered, and this one will suffice for now.

Towards the end of Friday, a few of the subsystems were checked out.  The Dewar pumped down fine, and the shutter appears to operate correctly.  The ISP computer came up and is now monitoring the many temperature sensors within the instrument.  More next week.

Todd



Thursday, July 26, 2012

The Installation details

 You only begin to get an idea of the scale of ODI when you see the pieces being assembled.  Here are a few photos from the installation procedure.  ODI goes on the telescope in 6 pieces - plus all the connections and peripheral systems like power supplies and chillers.  The first part, a shim (1) (machinable if we have to adjust the position or tilt of the overall instrument - ugh) attached to the rotator, was already on when we started on Wednesday. 

The next part is the forward corrector (2).  Here you see Brent and Mark guiding it into the port, where its flange will seat against the shim.  This piece weighs about 300 pounds.  All the optics have covers on them (mostly nice metal plates with handles and hardware for attaching them, but a few are just cardboard disks).  The covers are only removed at the very last minute to avoid damage or dust getting on the surface of the optics.










The next piece to go on is the ISP (3), which holds filter arms and ADC prisms.  This is mechanically the most complex part of the instrument with many motors and the electronics and wiring that drive them, gears, and sensors to monitor the positions of everything.  The ISP has almost all the moving parts in the system.  Here Mark and Brent guide the ISP over to the front corrector, already on the telescope.  The ISP weighs a little over 1000 pounds.  All the lifting was done with the WIYN dome crane, with the addition of an attachment called a Hydra Set Precision Load Positioner.  This hangs between the crane hook and the item being lifted and allows very precise (a few thousandths of an inch) vertical adjustment.  The green strap seen in this photo connects the Hydra Set to the lifting fixture attached to the top of the ISP.



This photo shows Charles and Brent guiding the ISP against the flange of the forward corrector.  The front cover of the ISP, which includes the shutter, has been removed because the bolts to attach it to the forward corrector are only accessible this way.   You can see the hubs for the three sets of filter arms, and the rearmost filter arm in each group.  The filters are held in metal frames that attach to the ends of these arms.  The filter removal and installation process involves removing the outer covers external to each filter arm group (the "chevron-shaped" pieces just counter-clockwise from each hub), running an arm all the way out, and then unclipping or clipping on the filter holder frame.
 
Now the cover (4) has been attached to the ISP.  The shutter is the rectangular black piece behind the cardboard disk.  At this point we had to do some work on the internal wiring.  One of the modifications necessary during the final lab testing was a redesign and rebuild of some of the worm-gear sets that run the filter arms in order to encase them in grease.  This resulted in our having to change the route of some of the wiring in such a way that we couldn't get the cover housings over the motor-gear assemblies.  Shelby made some cable extenders so that we could route them a different way that allowed us to put the covers on.  Charles here is removing one of those covers so we can add the extenders and connect the cables.  You can also see the hanger above the instrument that guides the wires, cables and hoses between ODI and the telescope.

Also hanging on the telescope is the computer (5) that controls the ISP.  This was assembled and programmed by Shelby Gott, whose back is seen here, while Daniel demonstrates his lifting abilities.  The computer is being attached to a support frame that is behind it.








The last piece to be installed is the Dewar (6).  It weighs about 1200 pounds (if I remember correctly).  It has the focal plane, the vacuum and cooling systems, and the Stargrasp controllers (the two gold boxes on top and bottom).  The Dewar is held by a lifting fixture that keeps it from tilting as it meets the ISP.   The filter arm hubs fit through holes in the flange of the Dewar.  Here Charles, Mark, and Brent (on the far side) guide the Dewar up to the ISP.










In this last picture, Mark and Brent are tightening the bolts that hold the Dewar to the ISP. All the bolts that hold the major parts of the instrument together are tightened with a torque wrench so that they are held securely, but not overtightened.








After finishing up the installation, we ran the field rotator around a few times to make sure that everything would clear and the balance was OK.  No problems.  The entire installation sequence pictured above took about 4 hours.  Mark will spend the next four days completing all the connections.  He's done this a couple of times in the lab, but there will be a few differences on the telescope.  If all goes well, we will begin to turn things on around the middle of next week.

Todd

Wednesday, July 25, 2012

ODI shipping, lifting time lapse movie

Pete Marenfeld took some very nice images and time-lapse videos of the ODI lifting yesterday. Enjoy a half day of work compressed into a few seconds. We hope that we can show more of these movies.

Daniel

[Edit: Unfortunately,  blogger.com compresses the videos almost beyond recognition. We will try to host high-resolution versions somewhere else.]



video

video
On Mission Road

Approaching Kitt Peak

Almost there



We bolted ODI on, and it didn't fall off…


… that would be the overly abridged summary of the day. In more detail: Before lunch the forward corrector was installed to the instrument rotator, directly followed by the instrument support package. After lunch we proceeded to install the dewar, which consumed the rest of the afternoon. At the end of the day we confirmed that the instrument rotator operated without any problems or interference. 

Shelby Gott installed the ODI instrument control computer to the telescope fork. Overall a busy day without surprises. 

Tomorrow we will start to establish all the electric, fluid, and gaseous connections. 

Daniel

ODI installed at the telescope.

Tuesday, July 24, 2012

Another hauling day

Today was a long and eventful (in the most positive sense) day. As planned, the ODI truck left the NOAO facilities around 6 am, arriving at Kitt Peak shortly before 8 am. By 10 am the two crates carrying the dewar and instrument support package were staged on the ground floor at WIYN. The shock loggers confirm a very smooth ride again – thanks to a very diligent driver and Gary Muller's well designed crates.

Before lunch both crates were unboxed and both ODI parts lifted to the dome floor. There we verified all the moving parts of ODI (they move!) and inspected the two prisms in the instrument. We conclude that there is no transportation damage evident, and ODI has been cleared for installation tomorrow. 

Finally, we got the ODI data acquisition computer installed in WIYN's computer cabinet, and Andrey has started to integrate it into the local network.

Daniel

Pictures below are partly by Heidi Schweiker. Pete Marenfeld took some more professional pictures that we hope to post soon. 


The truck approaches Kitt Peak

Charles Corson lifts the ODI dewar in the WIYN dome.

Andrey Yeatts installs the ODI data acquisition computer.


Monday, July 23, 2012

Loading ISP and Dewar for trip up to the telescope

Every day is a milestone now.  The "big" truck showed up at about 2:45 PM, and by 3:15 the two big crates were loaded.  An hour later, they were covered with tarps and ready to go.  The plan is to leave the truck in the NOAO yard overnight and drive up to the mountain at 6 AM tomorrow.  We were a bit worried about the boxes sitting in the heat for the next few hours (temp ~ 100 now), but the sky is getting cloudy and I'm hearing a little thunder.  Hope those tarps are waterproof.

Tomorrow, we will unload the crates at the WIYN dome, and open them up to make sure there is no visible damage.  Accelerometers will be removed and read out.  Then, we'll lift the ISP up to the dome floor, and hook it up to its control system and confirm that it is operational.  The rest of the day will be spent opening boxes and documenting the unpacking process.  If all goes well, we will start installation on Wednesday.

Todd





Friday, July 20, 2012

Shipping Day 1

The WIYN telescope ready for ODI
After a smooth ride (the shock logger on the forward corrector registered a maximum load of 0.75g), the NOAO cargo truck carrying the precious ODI parts arrived at 10 am at Kitt Peak. While waiting for the truck to arrive we took a sneak look at the telescope. The WIYN port is empty and ready for pODI with all the supply lines (including the new 30Amp power line). 


Once the truck arrived we wasted no time to start unloading, and it took us less than an hour to move all the crates and boxes inside the WIYN building. This completed our primary mission of the day. 


But why stop there? We used the time to set up the new ODI observing computer (2x30 inch displays, why would you settle for  less?), installed the computer rack in the computer cabinet, and started fit-checking and testing some equipment. As we speak, Gary and Charles are installing the interface shim (a giant washer) to the instrument rotator on the telescope.

Finally, a little cricket seems to be in the control room, cricketing along, and provides is very own soundtrack.


Daniel

Unloading of the truck.
Gary, Charles, and Heidi demonstrate
the ODI lifting fixture

Thursday, July 19, 2012

Ready for Trip #1

Most of the day was spent finishing packing and loading the NOAO cargo truck. Bill Porter could thread a needle with a forklift, and, with his help, we got the truck loaded up with everything (other than the two big crates - the ISP and the Dewar) that is too big to ride in the back of a van.  Finishing the filter holders and their shipping crate became a significant effort along the way.  Lots of people helped, and we're ready to go tomorrow morning.

Todd
Roger Repp working on a filter holder

Gary Poczulp working on a filter holder

Bill Porter maneuvering the forward corrector crate onto the truck while Gary and Daniel supervise.

Wednesday, July 18, 2012

Second day of packing

Got the ISP crate assembled today (see photo below), and tested Gary's lifting fixture for the crates.  Much of the smaller stuff is ready to go - some in containers.  We had a meeting to develop a good procedure for installing the filter modules on the arms (see other photo below).  Tomorrow, we load most of the smaller stuff onto the KPNO cargo truck with the forward corrector crate.  It will go up Friday at 8 AM and will be unloaded in the WIYN dome.  Monday we will finish crating up the Dewar, and the big truck will show up in the afternoon for loading before next Tuesday's early morning trip up.

Todd

Plastic wrapped ISP on its cart as crate is assembled around it,.  (Photo by Mark)
Gary, Daniel, and Chuck Gessner working out procedure for installing filter holders

Tuesday, July 17, 2012

First day of packing

We spent a lot of the day getting organized.  Translation: things didn't go as quickly as I hoped.  But, all in all, we made some progress and we are in good shape for our two transport days -Friday and Tuesday.  The ISP is almost ready.  It's on its shipping cart, all shrouds and covers are installed, and the shutter is attached.  Tomorrow we will finish the ISP crate and get the forward corrector crate done.  We also will try to get the myriad of smaller items - cables, hoses, bolts, tools, power supplies.... ready to go.  Picture below is Mark guiding the ISP onto it's shipping cart.  You can see the springs under the legs of the cart to protect it from shocks.  We will attach accelerometers to each of the three big pieces to make sure that they are handled gently.

Todd


The last days in the Assembly Hall

Following up on yesterday's post on the Instrument Support Package, the ODI dewar is also getting ready for shipping. The picture below shows the dewar on its shipping crate (yellow), awaiting the final tweaks before the great transport next week. Visibly attached to the dewar  are the large shutter (in black, made by the University of Bonn), and the two golden boxes that contain the Stargrasp CCD controller system (made by University of Hawaii).

Hidden on the left side you can spot an old-fashioned overhead projector. We used this ancient piece of technology to actually illuminate the large focal plate.

ODI Dewar in NOAO's assembly hall.

Monday, July 16, 2012

ISP ready for its shipping crate

This is the last crane-lift of the ISP on its cart until it is in its shipping crate.  We lifted it out of the environmental chamber, where filter arms and ADC prisms had been tested down to 0 C.  It is now sitting in the flex rig, where Gary Poczulp is cleaning the optics.  It will be crated up in the next few days, and it will travel up to the telescope on Tuesday, July 24.

Todd

No, it's not the spaghetti monster

This is the inside of the hex panel, which sits behind the focal plane.  This is where we hide all the connections and reorganization of signals, power, and cooling liquid.  The outside has a nice, orderly arrangement of connectors, all labeled.  One of Mark's last tasks before packing was to repair a small glycol leak in here.
 
Todd (photo by Daniel)

Friday, July 13, 2012

Readiness Review is behind us

Yesterday afternoon we passed our Readiness and Safety Review - the last hurdle before we begin packing up pODI for shipment to the mountain.  Mark, Daniel, Gary M, Charles, and Andrey assisted me in presenting the information and answering questions from the review panel.  The panel made about a dozen recommendations; almost all of them are suggestions for improving our procedures.  The panel was unanimously complimentary about the work that the team has done, and the detail in which all the steps have been thought through.  At the conclusion of the review, the relevant Directors, Pat for WIYN, Lori Allen representing Tim Beers for Kitt Peak, and David Sprayberry for downtown engineering services, all agreed that we are ready to go.

Mark is now doing a few last things to the Dewar, and we will begin packing next Tuesday.

Todd

Tuesday, July 10, 2012

Detector testing has stopped

Today at 4:30pm we have stopped active detector testing, and the next time we will power up the devices again will be on the mountain. Right now the dewar is on the vacuum pump and warming up.

We used the last hours of testing time to acquire guide star videos, which fundamentally worked.  The coming days will be used to harmonize the guide star data flow on the test dewar.


Beginning the transition to installation and commissioning

Today we stop testing (primarily detector testing) and begin the transition from development to installation and commissioning.  The material for the readiness review (to be held the day after tomorrow) has been distributed to the review panel and the directors who must sign off that we are ready to begin installation.  We hold our last weekly meeting of the ODI team tomorrow morning.  I don't see anything at this point to prevent our moving forward (fingers crossed).

Mark has about three days to finish up work on the Dewar, and we begin packing next Tuesday.  Daniel, Gary M, and I have to meet soon to finalize our guidelines for the CCD Mosaic filter inserts.

Todd

Friday, July 6, 2012

The pODI installation/commissioning calendar

Below you will see the calendar for the 2012B semester (plus a little more) showing the pODI activities from the readiness and safety review (July 12) through shipping, installation, engineering verification and commissioning.  The dates colored yellow on the calendar show when other instruments will be in use.  The dates colored blue have been set aside for preliminary shared risk observations by the WIYN partner scientists who have helped with commissioning.  Regular shared risk will start in March 2013.

Todd

Thursday, July 5, 2012

Status update to Board and SSC

Report on pODI Status - July 5, 2012

We are approaching the end of laboratory integration and testing of pODI, and beginning to prepare for shipping, installation, and commissioning.  We have encountered the normal number of problems that only become evident when you reach this point in the development of an instrument as complex as this.  So far, no show stoppers.  Here is a summary.

All 13 of the detectors in the pODI focal plane are operating properly, as are the focus sensors.  We have run the entire focal plane in "static imaging" mode, in which the central 3 X 3 array (approximately 24 arcmin square) integrates with amps turned off while the outer 4 are left active to acquire guide star data.  There is no detectable leakage of the glows from the outer, active detectors to the central region.  We have measured read noise, full well capacity, dark current, linearity and these all seem normal.  We are able to read small regions around a point representing a guide star at 20 Hz.  We have not yet run the detectors in orthogonal transfer (OT) mode on the focal plane (though we did in the test Dewar).  This is primarily due to a firmware upgrade that IfA put into the Stargrasp controllers that "broke" this mode. We'll get a fix from them in the next few days.  The IfA Stargrasp team visited and helped us with the optimization of the detectors.  We ended up with a small punch list of remaining work that they will do before we close out the contract. 

The Dewar has performed well in holding vacuum and maintaining operating temperature.  There is a little work to do on connectors, which will take place after we complete the detector testing.

The ISP, which holds the filter arms and ADC prisms, has been tested apart from the Dewar.  We had a significant problem with the filter arms, which are driven by two sets of worms and gears, a small, high-speed set and a large, low-speed set.  We had to modify the small ones by encasing them in a gearbox filled with grease, and add shims between filter arms to position the large ones more correctly.  We have now completed testing in our environmental chamber down to 0 degrees centigrade, and all functions seem to work acceptably.

Software is a major effort at this point.  On the instrument side, work is divided between supporting the detector testing and developing the capabilities that we will need for the commissioning effort.  The user interface is starting to take shape.  On the pipeline, portal, and archive side, the team is beginning acceptance testing of the first set of functional pieces.  Not all the pieces are there yet, so I won't call it an integration test, but we are trying to develop a model for formal testing of the complete system.  There will be an integrated system to test at the end of August with all the functionality needed to support commissioning.

Work on the telescope to prepare for ODI installation is also proceeding.  Delays in the dome repair work and aluminization have caused us to postpone our shipping date by a few days, but we have enough contingency in the installation schedule that this should not impact our overall schedule.  Our current plan is:

July 10 - completion of testing
July 12 - readiness and safety review
July 17 - begin packing
July 20 - first trip to mountain (front corrector and assorted small items)
July 24 - second trip to mountain (two large crates: Dewar and ISP)
July 27 - begin installation
Aug 16 - start engineering verification

Although the nominal date for first opening the dome is August 16, we will open earlier if things go well, and if the weather makes it possible.  Stay tuned!

Todd

Why a blog?

Today, in our weekly ODI team meeting, I said that next week would be our last meeting.  We will be getting ready to pack up ODI, ship it to the mountain, install it on the telescope, and begin commissioning activities.  I said that at that time, we would begin daily briefings to track status and to plan the upcoming work.  My idea had been to record those daily briefings on Google Docs, but Dick Shaw suggested that we try a Google blog instead.  That would allow interested people in the WIYN community to subscribe to the blog and keep up with our progress in a fairly simple way.

So, we will try this.  We (Todd or Daniel to start) will record items of interest in this blog.  After a short trial run with a limited readership, we will announce the blog to the Board and SSC, and to other appropriate groups.

Todd